Consider an RLC circuit consisting of a charged 500-
Answer #1

in this problem we need to convert time domain to frequency domain

transfer function=I(s)/V(s)

I(s)/V(s) =CS/(LCS2+RCS+1)

characteristic equation

LCS2+RCS+1=0

1)

R=0

W=sqrt(1/LC)

=sqrt(1/3*10-3*500*10-6)

=sqrt(6.66*105)

=0.81*103

f=0.81*1000/2*3.14

=129.36Hz

b)

W=R/2L

=16/2*32*10-3

=250

F=250/2*3.14

=39.8Hz

c)

for rhis we will get complex conjugate roots

d)

here we will get two different frequencies according to roots of characteristic equation